OpenStudy (anonymous):

Solve the following equation. 9x -3.2 = 6.8 + 8x

OpenStudy (anonymous):

9x-8x = 6.8+3.2 x = 6.8+3.2 x = _________

OpenStudy (saurabh135):

10